OpenStudy (anonymous):

limit as x approaches 0 (1-cos^2x/x)

OpenStudy (solomonzelman):

\[\lim_{x \rightarrow 0}\frac{1-\cos^2x}{0}\]

OpenStudy (anonymous):

its over x not 0

OpenStudy (solomonzelman):

\[\lim_{x \rightarrow 0}\frac{1-\cos^2x}{x}\]

OpenStudy (anonymous):

right

OpenStudy (solomonzelman):

Have you learned the L'Hopital's rule?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

L'Hospital's *

OpenStudy (solomonzelman):

Ohh, okay.

OpenStudy (anonymous):

what is it?

OpenStudy (solomonzelman):

it is a rule, that when a limit is indeterminate form (after plugging the value that x approaches), THEN you can take the derivative on top and bottom.

OpenStudy (solomonzelman):

But if you haven;t learned about derivatives yet, then don't worry about it.

OpenStudy (solomonzelman):

have you learned derivatives though?

OpenStudy (anonymous):

yes i know that the bottom is 1

OpenStudy (anonymous):

but im having trouble finding the derivative of the top

OpenStudy (anonymous):

the

OpenStudy (anonymous):

 part

OpenStudy (anonymous):

which is equal to  right?